Louisiana State University Health Sciences Center Administration & Finance
 

Tuesday, September 12, 2006

  • Dental School:
    • Pipe Insulation labeling/identification continues. The contractors have completed the labeling/identification of the ice water pipe and are continuing to label the remaining water pipes (Domestic Hot & Cold) in the basement. The pipe insulators are also changing some insulation that was damaged by construction.
    • Soft Water System piping is complete in the Power Plant. The contractor filled the tanks & added the salt mixture. All that remains is the start-up & testing of the system. Start-up will commence once the repair of the Motor Control Center has completed. This change order is being reviewed by the Architect.
    • Mold Remediation of heavily molded rooms (6413B, 6409, & 7409A) is complete. The contractors have removed the safety barriers around these areas & are awaiting inspection of the work.
    • The work remaining on the Emergency Generator is the swapping of connections from the temporary-temporary to the temporary-permanent.
    • The damaged vehicles stored at the Dental School Campus have been removed.
  • MEB:
    • Large Lecture Rooms. The furniture installer finished installing the electrical components today and will finish installing the race way covers when the pedestals come in for the ADA seating in the last row on each end of the projection booth. The replacement tops are scheduled to come in on the 19th.
    • The electricians are needed to connect the last whips to the power at the 3 of the desks.
  • 1542 Tulane Ave:
    • Entergy came out today and together with Ordes began pulling wires into the vault to supply power to all of the new panels in the basement. Equipment will have to be changed out in the vault before power can be restored. The equipment for the temporary power is on order and should be here by Thursday.
    • The cables from the vault to the 13.8 Switchgear are being connected at the switchgear.
    • They have begun connecting the cables to the power and the feeds to the 4th floor riser. The grounds are also being run and connected.
    • Labeling of the new panel boxes continues that Ordes have finished connecting to the transformer and the equipment they will feed.
    • Wink Engineering is working on plans to provide power and replace mechanical components that will make the building more suitable for occupancy.
  • Central Utility Plant:
    • The VFD is scheduled to arrive Monday of next week. They will mobilize next week to start work again. The line filters have arrived and are stored at the electricians warehouse. The motor is still not ready.